1. Two point charges q1 = 5.25 x 10-9 C and q2 = -8.75 x 10-9 C are 1.00 m apart in air. (a) What



force exist between them? Is this an attractive or a repulsive force? (b) The charges are allowed



to touch and separate again by 1.00 m. What force exists between them now? Is this an



attractive or a repulsive force?

1
Expert's answer
2022-03-09T10:32:26-0500

The Coulomb's law says


"F=k\frac{|q_1||q_2|}{r^2}"

(a)


"F=9*10^9*\frac{5.25*10^{-9}*8.75*10^{-9}}{1.00^2}=4.13*10^{-7}\:\rm N"


The force is attraction.

(b)


"q_1'=q_2'=\frac{q_1+q_2}{2}=-3.5*10^{-9}\:\rm C""F=9*10^9*\frac{3.5*10^{-9}*3.5*10^{-9}}{1.00^2}=1.10*10^{-7}\:\rm N"

The force is repulsion.